# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 3.2. The absolute movable cone conjecture
By [undefat, Lemma 6.3.(3), Theorem 6.4.(3)], we have $\Gamma\simeq\rho(W_{X})\rtimes\Gamma_{\mathop{\rm Bir}\nolimits}$. Since $T(X)_{\mathop{\rm tf}\nolimits}$ is the orthogonal complement of $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ in $H^{2}(X,\mathbb{Z})_{\mathop{\rm tf}\nolimits}$, elements of $W_{X}$ act trivially on $T(X)_{\mathop{\rm tf}\nolimits}$, therefore $\Gamma_{T}\simeq\rho(W_{X})\rtimes\Gamma_{T,\mathop{\rm Bir}\nolimits}$. Consequently, $\Gamma_{T,\mathop{\rm Bir}\nolimits}$ is finitely generated as a quotient of the finitely generated group $\Gamma_{T}$. By [undefat, Lemma 6.3.(2) and (3), Theorem 6.4.(3)], we have $\ker(\rho)\subset\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Bir}}\nolimits(X)$, where $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Bir}}\nolimits(X)$ is the subgroup of $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits$ of locally trivial monodromy operators induced by birational transformations of $X$, and moreover $\Gamma_{\mathop{\rm Bir}\nolimits}\simeq\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Bir}}\nolimits(X)/\ker(\rho)$. This gives us the following identifications
@@
so that $\left\{f\in\mathop{\rm Bir}\nolimits(X)\penalty 10000\ \middle|\penalty 10000\ f^{*}|_{H^{2}(X,\mathbb{Z})}=\text{id}\right\}\subset\widetilde{G}$. If we show that $\widetilde{G}$ is finite, we are done. Let $X\hookrightarrow\mathbb{P}^{N}$ be the embedding defined by $\left\lvert H\right\rvert$, where $N=\dim\left\lvert H\right\rvert$. Denote by $[X]$ the Zariski-closed point of $\mathop{\rm Hilb}\nolimits_{\mathbb{P}^{N}}$ corresponding to the embedding we have just written. It follows that $\widetilde{G}\subset\mathop{\rm PGL}\nolimits(N)$ and that $\widetilde{G}$ is the stabilizer of $[X]$ under the natural action of $\mathop{\rm PGL}\nolimits(N)$ on $\mathop{\rm Hilb}\nolimits_{\mathbb{P}^{N}}$. Since this action is Zariski-continuous, it follows that $\widetilde{G}$ is -Zariski-closed in $\mathop{\rm PGL}\nolimits(N)$. But now $\mathop{\rm PGL}\nolimits(N)$ is affine noetherian, so is $\widetilde{G}$, and since $\dim\mathop{\rm Aut}\nolimits(X)=h^{0}(X,{\mathcal{T}}_{X})=0$ by [undefd, Lemma 4.6], we have $\dim\widetilde{G}=0$ as well. It follows that $\widetilde{G}$ is finite, and that completes the proof. ∎
@@
Let $F$ be a field of characteristic $0$ and let $X$ be a projective primitive symplectic variety with $\mathbb{Q}$-factorial terminal singularities and $b_{2}(X)\geq 5$. The image of $\rho$ is a finite index subgroup and the kernel of $\rho$ is a finite subgroup of $\mathop{\rm Aut}\nolimits(X)\subset\mathop{\rm Bir}\nolimits(X)$.
@@
First, assume that $F$ is algebraically closed. Fix a finitely generated subfield $F^{\prime}\subset F$ such that $X$ is the base change of a projective variety $X^{\prime}$ over $\overline{F^{\prime}}\subset F$. After fixing an embedding $\overline{F^{\prime}}\subset\mathbb{C}$, we denote by $X^{\prime}_{\mathbb{C}}$ the base change of $X^{\prime}$ to $\mathbb{C}$. Note that both $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ are projective primitive symplectic varieties with $\mathbb{Q}$-factorial terminal singularities thanks to Lemma 2.3, Lemma 2.4 and Lemma 3.8. Moreover, both $X$ and $X^{\prime}$ have a second Betti number greater than or equal to $2$ thanks to [undefaac, VI - Corollary 4.3]. Now observe that
@@
by Lemma 3.9 and that $R_{X}\simeq R_{X^{\prime}}\simeq R_{X^{\prime}_{\mathbb{C}}}$ since, in our case, prime exceptional divisors correspond bijectively by base change. Moreover, we also have
@@
thanks to Lemma 3.25. Therefore, if $F$ is algebraically closed, we see that it suffices to prove the proposition for $F=\mathbb{C}$. In this case, the image of $\rho$ is a finite index subgroup by [undefat, Theorem 5.12 and Lemma 6.3]. Concerning the kernel of $\rho$, recall that a birational morphism that fixes an ample line bundle under pullback extends to an automorphism, so that $\ker(\rho)\subset\mathop{\rm Aut}\nolimits(X)$. Moreover, the group scheme $G$ of automorphisms of a projective variety fixing a quasi-polarization is a linear algebraic group as shown by [undefn, Proposition 2.26], so is of finite type over $\mathbb{C}$. Since $h^{0}(X,{\mathcal{T}}_{X})=0$ by [undefd, Lemma 4.6], $G$ is of dimension $0$ and is therefore finite, so that $\ker\rho\subset G$ is also finite.
We now turn to the case of a non-algebraically closed field $F$. Denote by $\overline{X}$ the base change of $X$ to $\overline{F}$, and by $G_{F}$ the absolute Galois group of $F$. Let $\overline{\rho}:\mathop{\rm Bir}\nolimits(\overline{X})\ltimes R_{\overline{X}}\to O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$. By the previous case, the kernel of $\overline{\rho}$ is finite and its image is of finite index in $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$. Note that $\mathop{\rm Bir}\nolimits(\overline{X})^{G_{F}}=\mathop{\rm Bir}\nolimits(X)$. The action of $G_{F}$ on $\mathop{\rm Bir}\nolimits(\overline{X})$ factors through a finite quotient since $\mathop{\rm Bir}\nolimits(\overline{X})$ is finitely generated by Lemma 3.29. Similarly, as $\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits}$ is finitely generated, the actions of $G_{F}$ on $R_{\overline{X}}=W_{\overline{X}}$ and $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$ factors through a finite quotient. Therefore, we can apply [undefm] to $\overline{\rho}$ and it follows that the induced morphism
@@
has a finite kernel, and its image is of finite index. Denote by $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits},\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ the subgroup of $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$ stabilizing $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ as a set. By [undefm, Proposition 2.2.(2)], $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})^{G_{F}}$ is of finite index in $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits},\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$. Finally, by [undefm, Proposition 2.2.(1)], the morphism $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits},\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})\to O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ has a finite kernel, and its image is of finite index. We then check directly that the composite
@@
Thanks to the results of this section, we will now prove Theorem A.(1). The strategy consists of adapting the proofs of [undefaat, Theorem 4.1.4] and [undefm, Corollary 3.15].
@@
is a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathop{\overline{\rm Pos}}\nolimits^{+}(X)$. We show that $\Pi\subset\mathop{\rm Mov}\nolimits(X)^{+}$. Take $x\in\Pi$. By Proposition 3.19.5, it is sufficient to show that $q_{\overline{X}}(x,E)\geq 0$ for any $E$ in a Galois orbit $I$ of prime exceptional divisors on $\overline{X}$ such that $W_{I}$ is finite, where $\overline{X}$ is the base change of $X$ to an algebraic closure $\overline{F}$ of $F$. Indeed, since $\Pi$ is rational polyhedral, we will automatically obtain the inclusion in $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$. For such an orbit $I$, after replacing the sum $E_{I}$ of the elements in $I$ by a sufficiently divisible multiple, we may assume that $E_{I}$ is Cartier. The longest element $r_{I}$ of the Coxeter system $(W_{I},I)$ acts on $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ as the reflection in the class of $E_{I}$. Since $r_{I}\in\Gamma$ and $x\in\Pi$, we obtain $q_{\overline{X}}(r_{I}(x),y)\geq q_{X}(x,y)$, which is equivalent to $q_{\overline{X}}(x,E_{I})=\left\lvert I\right\rvert\cdot q_{\overline{X}}(x,E)\geq 0$ for any $E\in I$.
Denote by $\Gamma_{\mathop{\rm Bir}\nolimits}$ the image of the morphism $\mathop{\rm Bir}\nolimits(X)\to O(\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}})$. We show that $\Pi$ is a fundamental domain for the action of $\Gamma_{\mathop{\rm Bir}\nolimits}$ on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$. By the above discussion, for any $x\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$, we can find $f\in\mathop{\rm Bir}\nolimits(X)$ and $r\in R_{X}$ such that $rf^{*}(x)\in\Pi\subset\mathop{\overline{\rm Mov}}\nolimits(X)$ (here, $f^{*}$ is the image of $f$ in $\Gamma$). Note that Proposition 3.20 implies that $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$ is a fundamental domain for the action of $R_{X}$ on $\mathop{\rm Pos}\nolimits(X)^{+}$. Observe that $f^{*}(x)$ already lies in $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$, therefore $f^{*}(x)=rf^{*}(x)\in\Pi$, and that concludes the proof. ∎
@@ 3.3. The absolute nef cone conjecture
Let $F$ be a field of characteristic $0$ and let $X$ be a projective primitive symplectic variety over $F$. If $X_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities, we define the birational ample cone of $X$ as the subset of $\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ defined by
@@
where $f$ runs through every birational maps with target a projective primitive symplectic variety $Y$ over $F$ such that $Y_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities.
@@
When $F=\mathbb{C}$, we recover the classical definition of the birational ample cone as written, for example, in [undefat, Definition 4.5]. Also, note that if $f:X\dashrightarrow Y$ is a birational map as in the definition above, then it is an isomorphism in codimension $1$ and the pullback of line bundles is therefore well defined. Indeed, the singularities of $X$ and $Y$ are terminal if and only if the singularities of $X_{\overline{F}}$ and $Y_{\overline{F}}$ are terminal by [undefao, Proposition 2.15], and then we can apply [undefap, Corollary 3.54].
@@
Let $F$ be a a field of characteristic $0$ and let $\overline{F}$ be an algebraic closure. If $X$ is a projective primitive symplectic variety over $F$ such that $\overline{X}:=X_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities, then we have the equality $\mathop{\rm BAmp}\nolimits(X)=\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$.
@@
For the inclusion $\mathop{\rm BAmp}\nolimits(X)\subseteq\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$, consider a birational map $f:X\dashrightarrow Y$ as in Definition 3.31, and take $\alpha\in\mathop{\rm Amp}\nolimits(Y)$. Set $\overline{Y}:=Y_{\overline{F}}$. We write the Cartesian square associated with the base change of $f$ to $\overline{F}$, where the vertical arrows are the canonical projections.
@@
For the reverse inclusion $\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\subseteq\mathop{\rm BAmp}\nolimits(X)$, let $G_{F}$ be the absolute Galois group of $F$. Take $\overline{x}\in\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ and consider a birational map $\overline{f}:\overline{X}\dashrightarrow\overline{Y}$ as in Definition 3.31 such that $\overline{x}=\overline{f}^{*}(\overline{y})$, with $\overline{y}\in\mathop{\rm Amp}\nolimits(\overline{Y})$. We claim that if $\overline{f}$ is defined over $F$, i.e. there exists a birational map $f:X\dashrightarrow Y$ over $F$ such that $\overline{f}=f_{\overline{F}}$, then we have proven the inclusion. Indeed, if that is the case, then the fact that $\overline{x}$ is $G_{F}$-invariant and that $\overline{f}$ is $G_{F}$-equivariant (being defined on $F$) implies that $\overline{y}$ is $G_{F}$-invariant, therefore $\overline{y}\in\mathop{\rm Amp}\nolimits(Y)$ by Lemma 3.4 together with the last observation in Remark 3.6.
We must therefore show that $\overline{f}$ descends to a birational map over $F$ via Galois descent. For vocabulary and definitions, see for example [undefl, $\mathsection$ 6.2]. Since $X$ is actually defined over $F$, we have a canonical descent data: for any $\sigma\in G_{F}$, there exists a scheme isomorphism $\phi_{\sigma}:\overline{X}\to\overline{X}$ such that the following diagram is commutative ($\sigma^{*}$ simply being $\text{id}_{X}\times\mathop{\rm Spec}\nolimits(\sigma)$).
${\overline{X}}$${\overline{X}}$${\mathop{\rm Spec}\nolimits(\overline{F})}$${\mathop{\rm Spec}\nolimits(\overline{F})}$$\scriptstyle{\phi_{\sigma}}$$\scriptstyle{\sigma^{*}}$
In addition, we have $p_{X}\circ\phi_{\sigma}=p_{X}$. Define the birational map $\psi_{\sigma}:=\overline{f}\circ\phi_{\sigma}\circ\overline{f}^{-1}:\overline{Y}\dashrightarrow\overline{Y}$ for all $\sigma\in G_{F}$. We see that the $\psi_{\sigma}$ fit into a commutative diagram as above, and that they are isomorphisms in codimension one. If we can show that these birational applications extend to isomorphisms, then they will define a descent data for both $\overline{Y}$ and $\overline{f}$, and we are done. Fix $\sigma\in G_{F}$. Note that since $\overline{x}\in\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ and $p_{X}\circ\phi_{\sigma}=p_{X}$, we have $\phi_{\sigma}^{*}(\overline{x})=\overline{x}$. We then have the equality
$\psi_{\sigma}^{*}(\overline{y})=(\overline{f}^{-1})^{*}\phi_{\sigma}^{*}(\overline{x})=(\overline{f}^{-1})^{*}(\overline{x})=\overline{y}$
showing that $\psi_{\sigma}^{*}$ fixes an ample class, and therefore extends to an isomorphism. This concludes the proof. ∎

### v2

@@ 3.2. The absolute movable cone conjecture
By [undefas, Lemma 6.3.(3), Theorem 6.4.(3)], we have $\Gamma\simeq\rho(W_{X})\rtimes\Gamma_{\mathop{\rm Bir}\nolimits}$. Since $T(X)_{\mathop{\rm tf}\nolimits}$ is the orthogonal complement of $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ in $H^{2}(X,\mathbb{Z})_{\mathop{\rm tf}\nolimits}$, elements of $W_{X}$ act trivially on $T(X)_{\mathop{\rm tf}\nolimits}$, therefore $\Gamma_{T}\simeq\rho(W_{X})\rtimes\Gamma_{T,\mathop{\rm Bir}\nolimits}$. Consequently, $\Gamma_{T,\mathop{\rm Bir}\nolimits}$ is finitely generated as a quotient of the finitely generated group $\Gamma_{T}$. By [undefas, Lemma 6.3.(2) and (3), Theorem 6.4.(3)], we have $\ker(\rho)\subset\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Bir}}\nolimits(X)$, where $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Bir}}\nolimits(X)$ is the subgroup of $\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Hdg}}\nolimits$ of locally trivial monodromy operators induced by birational transformations of $X$, and moreover $\Gamma_{\mathop{\rm Bir}\nolimits}\simeq\mathop{\rm Mon}^{\mathop{\rm 2,lt}}_{\mathop{\rm Bir}}\nolimits(X)/\ker(\rho)$. This gives us the following identifications
@@
so that $\left\{f\in\mathop{\rm Bir}\nolimits(X)\penalty 10000\ \middle|\penalty 10000\ f^{*}|_{H^{2}(X,\mathbb{Z})}=\text{id}\right\}\subset\widetilde{G}$. If we show that $\widetilde{G}$ is finite, we are done. Let $X\hookrightarrow\mathbb{P}^{N}$ be the embedding defined by $\left\lvert H\right\rvert$, where $N=\dim\left\lvert H\right\rvert$. Denote by $[X]$ the Zariski-closed point of $\mathop{\rm Hilb}\nolimits_{\mathbb{P}^{N}}$ corresponding to the embedding we have just written. It follows that $\widetilde{G}\subset\mathop{\rm PGL}\nolimits(N)$ and that $\widetilde{G}$ is the stabilizer of $[X]$ under the natural action of $\mathop{\rm PGL}\nolimits(N)$ on $\mathop{\rm Hilb}\nolimits_{\mathbb{P}^{N}}$. Since this action is Zariski-continuous, it follows that $\widetilde{G}$ is Zariski-closed in $\mathop{\rm PGL}\nolimits(N)$. But now $\mathop{\rm PGL}\nolimits(N)$ is affine noetherian, so is $\widetilde{G}$, and since $\dim\mathop{\rm Aut}\nolimits(X)=h^{0}(X,{\mathcal{T}}_{X})=0$ by [undefd, Lemma 4.6], we have $\dim\widetilde{G}=0$ as well. It follows that $\widetilde{G}$ is finite, and that completes the proof. ∎
@@
Let $F$ be a field of characteristic $0$ and $X$ a projective primitive symplectic variety with $b_{2}(X)\geq 5$ such that $\overline{X}:=X_{\overline{F}}$ has $\mathbb{Q}$-factorial terminal singularities. The image of $\rho$ is a finite index subgroup and the kernel of $\rho$ is a finite subgroup of $\mathop{\rm Aut}\nolimits(X)\subset\mathop{\rm Bir}\nolimits(X)$.
@@
First, assume that $F$ is algebraically closed. Fix a finitely generated subfield $F^{\prime}\subset F$ such that $X$ is the base change of a projective variety $X^{\prime}$ over $\overline{F^{\prime}}\subset F$. After fixing an embedding $\overline{F^{\prime}}\subset\mathbb{C}$, we denote by $X^{\prime}_{\mathbb{C}}$ the base change of $X^{\prime}$ to $\mathbb{C}$. Note that both $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ are projective primitive symplectic varieties with $\mathbb{Q}$-factorial terminal singularities thanks to Lemma 2.3, Lemma 2.4 and Lemma 3.8. Moreover, $b_{2}(X^{\prime})=b_{2}(X^{\prime}_{\mathbb{C}})\geq 5$ thanks to [undefaac, VI - Corollary 4.3]. Now observe that
@@
by Lemma 3.9, and that $R_{X}\simeq R_{X^{\prime}}\simeq R_{X^{\prime}_{\mathbb{C}}}$ since, in our case, the projections $X\to X^{\prime}$ and $X^{\prime}_{\mathbb{C}}\to X^{\prime}$ induce bijections between the sets of prime exceptional divisors. Moreover, we also have
@@
thanks to Lemma 3.26. Therefore, we are reduced to the case $F=\mathbb{C}$. In this case, the image of $\rho$ is a finite index subgroup by [undefas, Theorem 5.12 and Lemma 6.3]. Concerning the kernel of $\rho$, recall that a birational morphism that fixes an ample line bundle under pullback extends to an automorphism, so that $\ker(\rho)\subset\mathop{\rm Aut}\nolimits(X)$. Moreover, the group scheme $G$ of automorphisms of a projective variety fixing a quasi-polarization is a linear algebraic group as shown by [undefm, Proposition 2.26], so is of finite type over $\mathbb{C}$. Since $h^{0}(X,{\mathcal{T}}_{X})=0$ by [undefd, Lemma 4.6], $G$ is of dimension zero and is therefore finite, so that $\ker\rho\subset G$ is also finite.
We now turn to the case of a non-algebraically closed field $F$. Denote by $G_{F}$ the absolute Galois group of $F$. Let $\overline{\rho}:\mathop{\rm Bir}\nolimits(\overline{X})\ltimes R_{\overline{X}}\to O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$. By the previous case, the kernel of $\overline{\rho}$ is finite and its image is of finite index in $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$. Note that $\mathop{\rm Bir}\nolimits(\overline{X})^{G_{F}}=\mathop{\rm Bir}\nolimits(X)$. The action of $G_{F}$ on $\mathop{\rm Bir}\nolimits(\overline{X})$ factors through a finite quotient since $\mathop{\rm Bir}\nolimits(\overline{X})$ is finitely generated by Lemma 3.30. Similarly, as $\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits}$ is finitely generated, the actions of $G_{F}$ on $R_{\overline{X}}=W_{\overline{X}}$ and $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$ factors through a finite quotient. Therefore, we can apply [undefl, Lemma 3.12] to $\overline{\rho}$ and it follows that the induced morphism
@@
has a finite kernel, and its image is of finite index. Denote by $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits},\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ the subgroup of $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$ stabilizing $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ as a set. By [undefl, Proposition 2.2.(2)], $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})^{G_{F}}$ is of finite index in $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits},\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$. Finally, by [undefl, Proposition 2.2.(1)], the morphism
$O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits},\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})\to O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$
has a finite kernel, and its image is of finite index. We then check directly that the composite
@@
Thanks to the results of this section, we will now prove Theorem A.(1). The strategy consists of adapting the proofs of [undefaat, Theorem 4.1.4] and [undefl, Corollary 3.15].
@@
is a rational polyhedral fundamental domain under the action of $\Gamma$ on $\mathop{\overline{\rm Pos}}\nolimits^{+}(X)$. We show that $\Pi\subset\mathop{\overline{\rm Mov}}\nolimits(X)^{+}$. Take $x\in\Pi$. By Proposition 3.19.5, it is sufficient to show that $q_{\overline{X}}(x,E)\geq 0$ for any $E$ in a Galois orbit $I$ of prime exceptional divisors on $\overline{X}$ such that $W_{I}$ is finite, where $\overline{X}$ is the base change of $X$ to an algebraic closure $\overline{F}$ of $F$. Indeed, since $\Pi$ is rational polyhedral, we will automatically obtain the inclusion in $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$. For such an orbit $I$, after replacing the sum $E_{I}$ of the elements in $I$ by a sufficiently divisible multiple, we may assume that $E_{I}$ is Cartier. The longest element $r_{I}$ of the Coxeter system $(W_{I},I)$ acts on $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ as the reflection in the class of $E_{I}$. Since $r_{I}\in\Gamma$ and $x\in\Pi$, we obtain $q_{\overline{X}}(r_{I}(x),y)\geq q_{X}(x,y)$, which is equivalent to $q_{\overline{X}}(x,E_{I})=\left\lvert I\right\rvert\cdot q_{\overline{X}}(x,E)\geq 0$ for any $E\in I$.
Denote by $\Gamma_{\mathop{\rm Bir}\nolimits}$ the image of the morphism $\mathop{\rm Bir}\nolimits(X)\to O(\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}})$. We show that $\Pi$ is a fundamental domain under the action of $\Gamma_{\mathop{\rm Bir}\nolimits}$ on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$. By the above discussion, for any $x\in\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$, we can find $f\in\mathop{\rm Bir}\nolimits(X)$ and $r\in R_{X}$ such that $rf^{*}(x)\in\Pi\subset\mathop{\overline{\rm Mov}}\nolimits(X)$. Note that Proposition 3.20 implies that $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$ is a fundamental domain under the action of $R_{X}$ on $\mathop{\rm Pos}\nolimits(X)^{+}$. Observe that $f^{*}(x)$ already lies in $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$, therefore $f^{*}(x)=rf^{*}(x)\in\Pi$, and that concludes the proof. ∎
@@ 3.3. The absolute nef cone conjecture
Let $F$ be a field of characteristic $0$ and $X$ a projective primitive symplectic variety over $F$. If $X_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities, we define the birational ample cone of $X$ as the subset of $\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ given by
@@
where $f$ runs through all birational maps with target a projective primitive symplectic variety $Y$ over $F$ such that $Y_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities.
@@
When $F=\mathbb{C}$, we recover the classical definition of the birational ample cone as written, for example, in [undefas, Definition 4.5]. Also, note that if $f:X\dashrightarrow Y$ is a birational map as in the definition above, then $f$ is an isomorphism in codimension one, and hence the pullback of line bundles is well defined. Indeed, the singularities of $X$ and $Y$ are terminal if and only if the singularities of $X_{\overline{F}}$ and $Y_{\overline{F}}$ are terminal by [undefan, Proposition 2.15]. Therefore, the claim follows from [undefao, Corollary 3.54].
@@
Let $F$ be a a field of characteristic $0$ and $X$ a projective primitive symplectic variety over $F$ such that $\overline{X}:=X_{\overline{F}}$ has $\mathbb{Q}$-factorial and terminal singularities. Then $\mathop{\rm BAmp}\nolimits(X)=\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$.
@@
For the inclusion $\mathop{\rm BAmp}\nolimits(X)\subseteq\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$, let $f:X\dashrightarrow Y$ be a birational map as in Definition 3.32, and take $\alpha\in\mathop{\rm Amp}\nolimits(Y)$. Set $\overline{Y}:=Y_{\overline{F}}$. Consider the Cartesian square associated with the base change of $f$ to $\overline{F}$, where the vertical arrows are the canonical projections.
@@
For the reverse inclusion $\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\subseteq\mathop{\rm BAmp}\nolimits(X)$, let $G_{F}$ be the absolute Galois group of $F$. Take $\overline{x}\in\mathop{\rm BAmp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ and let $\overline{f}:\overline{X}\dashrightarrow\overline{Y}$ be a birational map as in Definition 3.32 such that $\overline{x}=\overline{f}^{*}(\overline{y})$, with $\overline{y}\in\mathop{\rm Amp}\nolimits(\overline{Y})$. We claim that if there exists a birational map $f:X\dashrightarrow Y$ over $F$ such that $\overline{f}=f_{\overline{F}}$, then we have proven the inclusion. Indeed, since both $\overline{x}$ and $\overline{f}$ are $G_{F}$-equivariant (being defined on $F$), $\overline{y}$ is also $G_{F}$-invariant and it follows that $\overline{y}\in\mathop{\rm Amp}\nolimits(Y)$ by Lemma 3.4 and Remark 3.6.
We must therefore show that $\overline{f}$ descends to a birational map over $F$. For definitions and terminology on Galois descent, see [undefk, $\mathsection$ 6.2]. Since $X$ is defined over $F$, we have a canonical descent data: for any $\sigma\in G_{F}$, there exists an isomorphism of schemes $\phi_{\sigma}:\overline{X}\to\overline{X}$ such that the diagram
${\overline{X}}$${\overline{X}}$${\mathop{\rm Spec}\nolimits(\overline{F})}$${\mathop{\rm Spec}\nolimits(\overline{F})}$$\scriptstyle{\phi_{\sigma}}$$\scriptstyle{\sigma}$
is commutative, and $p_{X}\circ\phi_{\sigma}=p_{X}$. For all $\sigma\in G_{F}$, define the birational map
$\psi_{\sigma}:=\overline{f}\circ\phi_{\sigma}\circ\overline{f}^{-1}:\overline{Y}\dashrightarrow\overline{Y}.$
Observe that the $\psi_{\sigma}$ fit into a commutative diagram as above, and that they are isomorphisms in codimension one. If we can show that these birational applications extend to isomorphisms, then they will define a descent data for both $\overline{Y}$ and $\overline{f}$. Fix $\sigma\in G_{F}$. Since $\overline{x}\in\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ and $p_{X}\circ\phi_{\sigma}=p_{X}$, we have $\phi_{\sigma}^{*}(\overline{x})=\overline{x}$, so that
$\psi_{\sigma}^{*}(\overline{y})=(\overline{f}^{-1})^{*}\phi_{\sigma}^{*}(\overline{x})=(\overline{f}^{-1})^{*}(\overline{x})=\overline{y}.$
Hence each $\psi_{\sigma}^{*}$ fixes an ample class and therefore extends to an isomorphism. This defines a descent datum for $\overline{Y}$ and $\overline{f}$, concluding the proof. ∎